A 0.240 M solution of the salt NaA has pH = 8.40. Calculate Ka for the acid HA.
Answer #1

Lets see the reaction

conjugate base A- reacts with water as

A- + H2O \rightleftharpoons HA + OH-  

We have

pH = 8.40

And we know that pH + pOH = 14

So pOH = 14 - pH = 14 - 8.40 = 5.6

pOH = -log[OH-]

then

[OH-] = 10-pOH = 10-5.6 = 2.5  \times 10-6 M

then

[HA] = 2.5  \times 10-6 M

For reaction

A- + H2O \rightleftharpoons HA + OH-  

equilibrium constant Kb will be

Kb = [HA][OH-] / [A-]

= (2.5  \times 10-6 )(2.5  \times 10-6 ) / 0.24 M

= 2.6 \times 10-11

we know that

Ka \times Kb = 1.0 \times 10-14  

then

Ka = 1.0 \times 10-14 / Kb = 1.0 \times 10-14 / 2.6 \times 10-11

= 3.8 \times 10-4   

So our answer is 3.8 \times 10-4